# Autonomous Bootstrapping of Quantum Dot Devices

### v1

@@ III.1 Bootstrapping: performance analysis
We initialize bootstrapping runs to test the performance of the autonomous bootstrapping algorithm presented in Fig. 1(b). In each test run, we extract the results of 1D gate characterization as well as the final sensor QD gate voltages obtained using rb-measurements.
@@
The finger gate characterization identifies working gates, and extracts pinch-off and saturation voltages. The five panels in the left-most column in Fig. 2(a) identify on the device SEM the finger gates used to form and control qubit QDs: OBL, PL, B, PR, and OBR. The corresponding 1D gate traces are shown in Fig. 2(b) and the extracted pinch-off and saturation voltages averaged over the tests are shown in Fig. 2(c). The error bars in Fig. 2 (c) represent one standard deviation within the collected sample and provide a metric of algorithm and device stability. High device and algorithm stability manifest through the low standard deviations of the extracted features: for the pinch-off values range from 111We use a notation value(uncertainty) to express uncertainties, for example, . All uncertainties herein reflect the uncorrelated combination of single-standard deviation statistical and systematic uncertainties. for OBL to for B; saturation values range from for B to for OBL. A summary of statistics for all finger gates is presented in Table 1.
@@
The middle column in Fig. 2(a) highlights the SP gate on the SEM. An example 1D trace used to extract a pinch-off, saturation, and parking value is shown in Fig. 2(b). The average pinch-off and saturation values collected at the characterization stage, as well as the parking positions obtained by detecting the most sensitive position on the Coulomb peak in rf-reflectometry, are depicted in Fig. 2(c). The standard deviations for pinch-off and saturation values of the SP gate extracted from current measurements are found to be and , respectively (see Table 1 for details). The standard deviation for parking voltage at the steepest slope of the Coulomb peak was at . Since pinch-off values are more well-defined physically, with the slope being steeper than at saturation, we expected the standard deviation of saturation values to be higher than those of pinch-offs.
@@
The two panels in the right-most column in Fig. 2(a) show the pairs of gates used to form a sensor QD. Fig. 2(b) shows the analyzed rb-measurements and Fig. 2(c) shows the statistics for the parking voltages in (c). The parking voltages for the BB, SBT and SBB gates were obtained using rb-measurements, with the corresponding 1D traces shown in the right-most panel in Fig. 2(b). Parking voltage for BB gate is the result of detecting a pinch-off in rb-measurement taken in combination with OBR, see Fig. 2(c) for performance statistics. The OBR is the result of the succeeding fine-tuning, as described in Sec. II.3. Finally, the sensor barriers’ parking positions correspond to the location of the most sensitive Coulomb oscillation peak, see Fig. 2(c). In this case, the dependence on the oscillatory behavior resulted in a larger standard deviation of the parking voltages for SBT and SBB gates, at 59 and 88 , respectively, than for BB and OBR, which rely on the well-behaving pinch-off values.
@@
To test bootstrapping performance, we measure a large charge stability diagram in the plunger space, with to qualitatively confirm that the qubit QDs can be sensed with the sensor QD. The choice of comes from prior work [30]. Out of bootstrapping test runs, were unsuccessful at charge sensing the qubit QDs: had uncharacteristically high parking voltage for the OBL that provided an insufficient current pinch-off to form the left QD, and two remaining runs failed to detect the Coulomb oscillations in the rb-measurements for SBT vs SBB measurements, and thus no sensor QD could be formed.
@@ III.2 Full tuning: performance analysis
Once the bootstrapping algorithm was validated, the algorithm was expanded to include an active coarse-tuning feedback routine for calibrating the V, B, and V gates to bring the QD device to a DD state, see Fig. 1(b). The full tuning testing involves running the bootstrapping algorithm followed by the DSE-based coarse tuning module as a single protocol, with the DSE module relying exclusively on rf-reflectometry.
@@
Figure 3(a) shows an overview of a sample coarse-tuning run with active feedback in the plunger gate space. A close-up of the tuning path is shown in Fig. 3(b). In this example, the QD device is initialized in a configuration corresponding to the SDL state, with . Then, through an iterative application of the DSE module along with the navigation routine, the plunger gates’ configuration is adjusted to bring the QD devices to the target DD state. Figure 3(c) depicts six scans with pixels resolution taken along the tuning path, indicating an increased likelihood of the DD state as the algorithm navigates toward the DD state.
Out of full tuning algorithm runs terminated at a gate configuration with DD likelihood of the state vector surpassing the predefined threshold of , 2 were qualitatively assessed as false positives. All remaining test runs were successful. While human experts might have been able to recover a DD state from these failed positives, in the time spent recovering, the algorithm could possibly have tuned from scratch again. The probability distributions of the DD likelihood for the initial and the final state vectors are shown in Fig.3(d). The histogram in Fig.3(e) shows the number of iterations it took for this algorithm to find the target state. We find a slight correlation between the probability of DD in the initial scan and the number of iterations. In particular, while for test runs with the number of iterations goes up to , for , . The average time spent by the algorithm from starting bootstrapping until termination by successful tuning to DD state was , with spent on bootstrapping and about on tuning. This compares very well with human expert tuning which typically requires hours to reach charge sensing of double QDs [22].
@@ IV Summary and outlook
We propose and experimentally demonstrate a modular autonomous bootstrapping algorithm for bringing a GaAs spin qubit QD device from a completely untuned state to the DD global state configuration. The algorithm performs gate characterization and device diagnostics, ensures the formation of a qubit QD channel, and prepares the charge sensing QD for rf-reflectometry. The low initialization failure rate and high consistency of the saturation and pinch-off voltages confirm the algorithm’s stability. The bootstrapping protocol efficiently deploys a combination of measurement techniques such as regular 1D voltage sweeps and rb-measurements for the current sweeps and small and fast 2D sweeps using rf-reflectometry, improving over previous demonstrations in the field that used transport measurements. Moreover, the high success rate for tuning to DD global state configuration confirms the utility of the ML-based DSE module, trained exclusively using simulated data, for deployment across varying device architectures and materials.
@@
In the present work, the rf-reflectometry circuit (frequency, phase, position in the I-Q plane, demodulation parameters) was set to known values optimized manually; in the future, the rf-reflectometry optimization could be automated using simple feedback based on similar measurements to those presented here. A broad application of our modules to different materials systems (e.g., nanowires, germanium quantum wells) might help understand materials differences, disorders, and future platform-specific improvements.
Finally, this work focuses on GaAs devices, where a single layer of gates is sufficient to confine the QDs. This simplifies tuning compared to QDs defined by overlapping gate electrodes, which may add complexities due to cross-capacitance as well as a higher number of parameters that need to be adjusted. While preliminary testing of some of the bootstrapping algorithm components shows good performance when tested on an accumulation mode Si sample, full cross-platform compatibility will require the development of specialized data analysis tools.
@@ Acknowledgements.
AC and TR acknowledge support from the Inge Lehmann Program of the Independent Research Fund Denmark. This work received funding from the US Army Research Office (ARO) under Award No. W911NF-24-2-0043, the European Union’s Horizon 2020 Research and Innovation Programme under grant agreement 101017733 (QuantERA II), and through the Horizon Europe Framework Programme’s IGNITE project under grant agreement No. 101069515. We thank Evert van Nieuwenberg for the valuable discussions and supervision of Lara Lausen. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by NIST.References
- Loss and DiVincenzo [1998] D. Loss and D. P. DiVincenzo, Quantum computation with quantum dots, Phys. Rev. A 57, 120 (1998).
@@
- Vandersypen and Eriksson [2019] L. M. K. Vandersypen and M. A. Eriksson, Quantum computing with semiconductor spins, Phys. Today 72, 38 (2019).
@@
- Bavdaz et al. [2022] P. L. Bavdaz, H. G. J. Eenink, J. van Staveren, M. Lodari, C. G. Almudever, J. S. Clarke, F. Sebasatiano, M. Veldhorst, and G. Scappucci, A quantum dot crossbar array with sublinear scaling of interconnects at cryogenic temperature, npj Quantum Inf. 8, 1 (2022).
@@
- Boter et al. [2022] J. M. Boter, J. P. Dehollain, J. P. van Dijk, Y. Xu, T. Hensgens, R. Versluis, H. W. Naus, J. S. Clarke, M. Veldhorst, F. Sebastiano, and L. M. Vandersypen, Spiderweb array: A sparse spin-qubit array, Phys. Rev. Applied 18, 024053 (2022).
@@
- Ansaloni et al. [2020] F. Ansaloni, A. Chatterjee, H. Bohuslavskyi, B. Bertrand, L. Hutin, M. Vinet, and F. Kuemmeth, Single-electron operations in a foundry-fabricated array of quantum dots, Nat. Commun. 11, 6399 (2020).
@@
- Watson et al. [2018] T. F. Watson, S. G. J. Philips, E. Kawakami, D. R. Ward, P. Scarlino, M. Veldhorst, D. E. Savage, M. G. Lagally, M. Friesen, S. N. Coppersmith, M. A. Eriksson, and L. M. K. Vandersypen, A programmable two-qubit quantum processor in silicon, Nature 555, 633 (2018).
@@
- Xue et al. [2021] X. Xue, B. Patra, J. P. G. van Dijk, N. Samkharadze, S. Subramanian, A. Corna, B. Paquelet Wuetz, C. Jeon, F. Sheikh, E. Juarez-Hernandez, B. P. Esparza, H. Rampurawala, B. Carlton, S. Ravikumar, C. Nieva, S. Kim, H.-J. Lee, A. Sammak, G. Scappucci, M. Veldhorst, F. Sebastiano, M. Babaie, S. Pellerano, E. Charbon, and L. M. K. Vandersypen, Cmos-based cryogenic control of silicon quantum circuits, Nature 593, 205 (2021).
@@
- Zwerver et al. [2022] A. M. J. Zwerver, T. Krähenmann, T. F. Watson, L. Lampert, H. C. George, R. Pillarisetty, S. A. Bojarski, P. Amin, S. V. Amitonov, J. M. Boter, R. Caudillo, D. Correas-Serrano, J. P. Dehollain, G. Droulers, E. M. Henry, M. Kotlyar, R.vLodari, F. Lüthi, D. J. Michalak, B. K. Mueller, S. Neyens, J. Roberts, N. Samkharadze, G. Zheng, O. K. Zietz, G. Scappucci, M. Veldhorst, L. M. K. Vandersypen, and J. S. Clarke, Qubits made by advanced semiconductor manufacturing, Nat. Electron. 5, 184 (2022).
@@
- Noiri et al. [2022] A. Noiri, K. Takeda, T. Nakajima, T. Kobayashi, A. Sammak, G. Scappucci, and S. Tarucha, Fast universal quantum gate above the fault-tolerance threshold in silicon, Nature 601, 338 (2022).
@@
- Philips et al. [2022] S. G. J. Philips, M. T. Madzik, S. V. Amitonov, S. L. de Snoo, M. Russ, N. Kalhor, C. Volk, W. I. L. Lawrie, D. Brousse, L. Tryputen, B. P. Wuetz, A. Sammak, M. Veldhorst, G. Scappucci, and L. M. K. Vandersypen, Universal control of a six-qubit quantum processor in silicon, Nature 609, 919 (2022).
@@
- Weinstein et al. [2023] A. J. Weinstein, M. D. Reed, A. M. Jones, R. W. Andrews, D. Barnes, J. Z. Blumoff, L. E. Euliss, K. Eng, B. Fong, S. D. Ha, D. R. Hulbert, C. Jackson, M. Jura, T. E. Keating, J. Kerckhoff, A. A. Kiselev, J. Matten, G. Sabbir, A. Smith, J. Wright, M. T. Rakher, T. D. Ladd, and M. G. Borselli, Universal logic with encoded spin qubits in silicon, Nature 615, 817 (2023).
@@
- Takeda et al. [2022] K. Takeda, A. Noiri, T. Nakajima, T. Kobayashi, and S. Tarucha, Quantum error correction with silicon spin qubits, Nature 608, 682 (2022).
@@
- Madzik et al. [2022] M. T. Madzik, S. Asaad, A. Youssry, B. Joecker, K. M. Rudinger, E. Nielsen, K. C. Young, T. J. Proctor, A. D. Baczewski, A. Laucht, V. Schmitt, F. E. Hudson, K. M. Itoh, A. M. Jakob, B. C. Johnson, D. N. Jamieson, A. S. Dzurak, C. Ferrie, R. Blume-Kohout, and A. Morello, Precision tomography of a three-qubit donor quantum processor in silicon, Nature 601, 348 (2022).
@@
- Chatterjee et al. [2022] A. Chatterjee, F. Ansaloni, T. Rasmussen, B. Brovang, F. Fedele, H. Bohuslavskyi, O. Krause, and F. Kuemmeth, Autonomous estimation of high-dimensional coulomb diamonds from sparse measurements, Phys. Rev. Appl. 18, 064040 (2022).
@@
- Burkard et al. [2023] G. Burkard, T. D. Ladd, A. Pan, J. M. Nichol, and J. R. Petta, Semiconductor spin qubits, Rev. Mod. Phys. 95, 025003 (2023).
@@
- Zwolak and Taylor [2023] J. P. Zwolak and J. M. Taylor, Colloquium: Advances in automation of quantum dot devices control, Rev. Mod. Phys. 95, 011006 (2023).
@@
- Ziegler et al. [2022] J. Ziegler, T. McJunkin, E. S. Joseph, S. S. Kalantre, B. Harpt, D. E. Savage, M. G. Lagally, M. A. Eriksson, J. M. Taylor, and J. P. Zwolak, Toward robust autotuning of noisy quantum dot devices, Phys. Rev. Appl. 17, 024069 (2022).
@@
- Zwolak et al. [2020] J. P. Zwolak, T. McJunkin, S. S. Kalantre, J. Dodson, E. R. MacQuarrie, D. Savage, M. Lagally, S. Coppersmith, M. A. Eriksson, and J. M. Taylor, Autotuning of double-dot devices in situ with machine learning, Phys. Rev. Appl. 13, 034075 (2020).
@@
- Ziegler et al. [2023a] J. Ziegler, F. Luthi, M. Ramsey, F. Borjans, G. Zheng, and J. P. Zwolak, Tuning arrays with rays: Physics-informed tuning of quantum dot charge states, Phys. Rev. Appl. 20, 034067 (2023a).
@@
- Zwolak et al. [2021] J. P. Zwolak, T. McJunkin, S. S. Kalantre, S. F. Neyens, E. R. MacQuarrie, M. A. Eriksson, and J. M. Taylor, Ray-based framework for state identification in quantum dot devices, PRX Quantum 2, 020335 (2021).
@@
- Berritta et al. [2024] F. Berritta, T. Rasmussen, J. A. Krzywda, J. van der Heijden, F. Fedele, S. Fallahi, G. C. Gardner, M. J. Manfra, E. van Nieuwenburg, J. Danon, A. Chatterjee, and F. Kuemmeth, Real-time two-axis control of a spin qubit, Nat. Commun. 15, 1676 (2024).
@@
- Moon et al. [2020] H. Moon, D. T. Lennon, J. Kirkpatrick, N. M. van Esbroeck, L. C. Camenzind, L. Yu, F. Vigneau, D. M. Zumbühl, G. A. D. Briggs, M. A. Osborne, D. Sejdinovic, E. A. Laird, and N. Ares, Machine learning enables completely automatic tuning of a quantum device faster than human experts, Nat. Commun. 11, 4161 (2020).
@@
- Podd et al. [2010] G. J. Podd, S. J. Angus, D. A. Williams, and A. J. Ferguson, Charge sensing in intrinsic silicon quantum dot, Appl. Phys. Lett. 96, 082104 (2010).
@@
- Durrer et al. [2020] R. Durrer, B. Kratochwil, J. Koski, A. Landig, C. Reichl, W. Wegscheider, T. Ihn, and E. Greplova, Automated tuning of double quantum dots into specific charge states using neural networks, Phys. Rev. Appl. 13, 054019 (2020).
@@
- Ziegler et al. [2023b] J. Ziegler, F. Luthi, M. Ramsey, F. Borjans, G. Zheng, and J. P. Zwolak, Automated extraction of capacitive coupling for quantum dot systems, Phys. Rev. Appl. 19, 054077 (2023b).
@@
- Zwolak et al. [2018] J. P. Zwolak, S. S. Kalantre, X. Wu, S. Ragole, and J. M. Taylor, QFlow lite dataset: A machine-learning approach to the charge states in quantum dot experiments, PLoS ONE 13, e0205844 (2018).
@@
- Fedele et al. [2021] F. Fedele, A. Chatterjee, S. Fallahi, G. C. Gardner, M. J. Manfra, and F. Kuemmeth, Simultaneous operations in a two-dimensional array of singlet-triplet qubits., PRX Quantum 2, 040306 (2021).
@@
- Baart et al. [2016] T. A. Baart, P. T. Eendebak, C. Reichl, W. Wegscheider, and L. M. K. Vandersypen, Computer-automated tuning of semiconductor double quantum dots into the single-electron regime, Appl. Phys. Lett. 108, 213104 (2016).
@@
- Note [1] We use a notation value(uncertainty) to express uncertainties, for example, . All uncertainties herein reflect the uncorrelated combination of single-standard deviation statistical and systematic uncertainties.

### v2

@@ III.1 Bootstrapping: Performance analysis
We initialize bootstrapping runs to test the performance of the autonomous bootstrapping algorithm presented in Fig. 1(b). In each test run, we extract the results of 1D gate characterization as well as the final sensor-QD gate voltages obtained using rb measurements.
@@
The finger-gate characterization identifies working gates, and extracts pinch-off and saturation voltages. The five panels in the leftmost column in Fig. 2(a) identify on the device SEM the finger gates used to form and control qubit QDs: OBL, PL, B, PR, and OBR. The corresponding 1D gate traces are shown in Fig. 2(b) and the extracted pinch-off and saturation voltages averaged over the tests are shown in Fig. 2(c). The error bars in Fig. 2(c) represent one standard deviation within the collected sample and provide a metric of algorithm and device stability. High device and algorithm stability manifest through the low standard deviations of the extracted features: for the pinch-off, values range from 111We use a notation value(uncertainty) to express uncertainties, for example, . All uncertainties herein reflect the uncorrelated combination of single-standard deviation statistical and systematic uncertainties. for OBL to for B; saturation values range from for B to for OBL. A summary of statistics for all finger gates is presented in Table 1.
@@
The middle column in Fig. 2(a) highlights the SP gate on the SEM. An example of a 1D trace used to extract pinch-off and saturation voltages is shown in Fig. 2(b). The average pinch-off and saturation values collected at the characterization stage, as well as the parking positions obtained by detecting the most sensitive position on the highest-scoring [according to the Eq. (2)] Coulomb peak in rf reflectometry after the formation of the sensor QD and the localization of the DD region is concluded, are depicted in Fig. 2(c). The standard deviations for pinch-off and saturation values of the SP gate extracted from current measurements are found to be and , respectively (for details, see Table 1). The standard deviation for the SP position selected on the highest-scoring Coulomb oscillation peak and parking at its steepest slope was at . Since pinch-off values are more well-defined physically, with the slope being steeper than at saturation, we expected the standard deviation of saturation values to be higher than those of pinch-offs.
@@
The two panels in the rightmost column in Fig. 2(a) show the pairs of gates used to form a sensor QD. In Fig. 2(b), we show the analyzed rb measurements and in Fig. 2(c), we show the statistics for the parking voltages in (c). The parking voltages for the BB, SBT, and SBB gates have been obtained using rb measurements, with the corresponding 1D traces shown in the rightmost panel in Fig. 2(b). The parking voltage for the BB gate is the result of detecting a pinch-off in rb measurement taken in combination with OBR [for performance statistics, see Fig. 2(c)]. The OBR is the result of the succeeding fine tuning, as described in Sec. II.3. Finally, the parking positions of the sensor barriers correspond to the location of the most sensitive Coulomb oscillation peak [see Fig. 2(c)]. In this case, the dependence on the oscillatory behavior has resulted in a larger standard deviation of the parking voltages for the SBT and SBB gates, at 59 and 88 , respectively, than for BB and OBR, which rely on the well-behaving pinch-off values.
@@
To test the bootstrapping performance, we measure a large charge stability diagram in the plunger space, with to qualitatively confirm that the qubit QDs can be sensed with the sensor QD. The choice of comes from prior work [32]. Out of bootstrapping test runs, were unsuccessful at charge sensing the qubit QDs. Among these failed tuning attempts, had an uncharacteristically high parking voltage for the OBL that provided an insufficient current pinch-off to form the left QD, and two remaining runs failed to detect the Coulomb oscillations in the rb measurements for SBT versus SBB measurements, and thus no sensor QD could be formed. The discovered failure modes have provided an experimental heuristic for invoking the sensor-QD re-tuning from the initial device state that was implemented for the full tuning protocol discussed in Sec. III.2.
@@ III.2 Full tuning: performance analysis
Once the bootstrapping algorithm had been validated, the algorithm was expanded to include an active coarse-tuning feedback routine for calibrating the V, B, and V gates to bring the QD device to a DD state [see Fig. 1(b)]. The full tuning testing involves running the bootstrapping algorithm followed by the DSE-based coarse-tuning module as a single protocol, with the DSE module relying exclusively on rf reflectometry.
@@
In Fig. 3(a), we show an overview of a sample coarse-tuning run with active feedback in the plunger gate space. A close-up of the tuning path is shown in Fig. 3(b). In this example, the QD device is initialized in a configuration corresponding to the SDL state, with . Then, through an iterative application of the DSE module along with the navigation routine, the configuration of the plunger gates is adjusted to bring the QD devices to the target DD state. In Fig. 3(c), we depict six scans with -pixels resolution taken along the tuning path, indicating an increased likelihood of the DD state as the algorithm navigates toward the DD state.
Out of full tuning algorithm runs terminated at a gate configuration with the DD likelihood of the state vector surpassing the predefined threshold of , two were qualitatively assessed as false positives. All remaining test runs were successful. While human experts might have been able to recover a DD state from these failed positives, in the time spent recovering, the algorithm could possibly have tuned from scratch again. The probability distributions of the DD likelihood for the initial and the final state vectors are shown in Fig. 3(d). The histogram in Fig. 3(e) shows the number of iterations it took for this algorithm to find the target state. We find a slight correlation between the probability of DD in the initial scan and the number of iterations. In particular, while for test runs with the number of iterations goes up to , for , . The average time spent by the algorithm from starting bootstrapping until termination by successful tuning to the DD state was , with spent on bootstrapping and about on tuning. This compares very well with human expert tuning which, typically requires hours to reach charge sensing of double QDs [23].
@@ IV Summary and outlook
We propose and experimentally demonstrate a modular autonomous bootstrapping algorithm for bringing a GaAs spin-qubit QD device from a completely untuned state to the DD global-state configuration. The algorithm performs gate characterization and device diagnostics, ensures the formation of a qubit-QD channel, and prepares the charge sensing QD for rf reflectometry. The low initialization failure rate and high consistency of the saturation and pinch-off voltages confirm the stability of the algorithm. The bootstrapping protocol efficiently deploys a combination of measurement techniques such as regular 1D voltage sweeps and rb measurements for the current sweeps and small and fast 2D sweeps using rf reflectometry, improving over previous demonstrations in the field that have used transport measurements. Moreover, the high success rate for tuning to the DD global-state configuration confirms the utility of the ML-based DSE module, trained exclusively using simulated data, for deployment across varying device architectures and materials.
@@
In the present work, the rf-reflectometry circuit (frequency, phase, position in the I-Q plane, demodulation parameters) has been set to known values optimized manually; in the future, the rf-reflectometry optimization could be automated using simple feedback based on measurements that are similar to those presented here. Additionally, recent developments in high-speed and high-sensitivity real-time “video mode” data acquisition could be complemented by this protocol, allowing for more rapid sensor-QD tuning for charge sensing of the QD arrays, eliminating the need for human intervention. A broad application of our modules to different materials systems (e.g., nanowires, germanium quantum wells) might help us to understand differences between materials, disorders, and future platform-specific improvements.
Finally, this work focuses on GaAs devices, where a single layer of gates is sufficient to confine the QDs. This simplifies tuning compared to QDs defined by overlapping gate electrodes, which may add complexities due to cross-capacitance as well as a higher number of parameters that need to be adjusted. While preliminary testing of some of the bootstrapping algorithm components shows good performance when tested on an accumulation-mode Si sample, full cross-platform compatibility will require the development of specialized data-analysis tools.
@@ Acknowledgements.
A.C. and T.R. acknowledge support from the Inge Lehmann Program of the Independent Research Fund Denmark. This work received funding from the U.S. Army Research Office (ARO) under Award No. W911NF-24-2-0043, the European Union Horizon 2020 Research and Innovation Programme under Grant Agreement No. 101017733 (QuantERA II), and through the Horizon Europe Framework Programme’s Integrated Germanium Quantum Technology (IGNITE) project under Grant Agreement No. 101069515. We thank Evert van Nieuwenberg for the valuable discussions and supervision of Lara Lausen. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by the National Institute of Standards and Technology (NIST).References
- [1] D. Loss and D. P. DiVincenzo, Quantum computation with quantum dots, Phys. Rev. A 57, 120 (1998).
@@
- [2] L. M. K. Vandersypen and M. A. Eriksson, Quantum computing with semiconductor spins, Phys. Today 72, 38 (2019).
@@
- [3] P. L. Bavdaz, H. G. J. Eenink, J. van Staveren, M. Lodari, C. G. Almudever, J. S. Clarke, F. Sebasatiano, M. Veldhorst, and G. Scappucci, A quantum dot crossbar array with sublinear scaling of interconnects at cryogenic temperature, npj Quantum Inf. 8, 1 (2022).
@@
- [4] J. M. Boter, J. P. Dehollain, J. P. van Dijk, Y. Xu, T. Hensgens, R. Versluis, H. W. Naus, J. S. Clarke, M. Veldhorst, F. Sebastiano, and L. M. Vandersypen, Spiderweb array: A sparse spin-qubit array, Phys. Rev. Applied 18, 024053 (2022).
@@
- [5] F. Ansaloni, A. Chatterjee, H. Bohuslavskyi, B. Bertrand, L. Hutin, M. Vinet, and F. Kuemmeth, Single-electron operations in a foundry-fabricated array of quantum dots, Nat. Commun. 11, 6399 (2020).
@@
- [6] T. F. Watson, S. G. J. Philips, E. Kawakami, D. R. Ward, P. Scarlino, M. Veldhorst, D. E. Savage, M. G. Lagally, M. Friesen, S. N. Coppersmith, M. A. Eriksson, and L. M. K. Vandersypen, A programmable two-qubit quantum processor in silicon, Nature 555, 633 (2018).
@@
- [7] X. Xue, B. Patra, J. P. G. van Dijk, N. Samkharadze, S. Subramanian, A. Corna, B. Paquelet Wuetz, C. Jeon, F. Sheikh, E. Juarez-Hernandez, et al. Cmos-based cryogenic control of silicon quantum circuits, Nature 593, 205 (2021).
@@
- [8] A. M. J. Zwerver, T. Krähenmann, T. F. Watson, L. Lampert, H. C. George, R. Pillarisetty, S. A. Bojarski, P. Amin, S. V. Amitonov, J. M. Boter, et al. Qubits made by advanced semiconductor manufacturing, Nat. Electron. 5, 184 (2022).
@@
- [9] A. Noiri, K. Takeda, T. Nakajima, T. Kobayashi, A. Sammak, G. Scappucci, and S. Tarucha, Fast universal quantum gate above the fault-tolerance threshold in silicon, Nature 601, 338 (2022).
@@
- [10] S. G. J. Philips, M. T. Madzik, S. V. Amitonov, S. L. de Snoo, M. Russ, N. Kalhor, C. Volk, W. I. L. Lawrie, D. Brousse, L. Tryputen, B. P. Wuetz, A. Sammak, M. Veldhorst, G. Scappucci, and L. M. K. Vandersypen, Universal control of a six-qubit quantum processor in silicon, Nature 609, 919 (2022).
@@
- [11] A. J. Weinstein, M. D. Reed, A. M. Jones, R. W. Andrews, D. Barnes, J. Z. Blumoff, L. E. Euliss, K. Eng, B. Fong, S. D. Ha, et al. Universal logic with encoded spin qubits in silicon, Nature 615, 817 (2023).
@@
- [12] K. Takeda, A. Noiri, T. Nakajima, T. Kobayashi, and S. Tarucha, Quantum error correction with silicon spin qubits, Nature 608, 682 (2022).
@@
- [13] M. T. Madzik, S. Asaad, A. Youssry, B. Joecker, K. M. Rudinger, E. Nielsen, K. C. Young, T. J. Proctor, A. D. Baczewski, A. Laucht, et al. Precision tomography of a three-qubit donor quantum processor in silicon, Nature 601, 348 (2022).
@@
- [14] A. Chatterjee, F. Ansaloni, T. Rasmussen, B. Brovang, F. Fedele, H. Bohuslavskyi, O. Krause, and F. Kuemmeth, Autonomous estimation of high-dimensional coulomb diamonds from sparse measurements, Phys. Rev. Appl. 18, 064040 (2022).
@@
- [15] G. Burkard, T. D. Ladd, A. Pan, J. M. Nichol, and J. R. Petta, Semiconductor spin qubits, Rev. Mod. Phys. 95, 025003 (2023).
@@
- [16] J. P. Zwolak and J. M. Taylor, Colloquium: Advances in automation of quantum dot devices control, Rev. Mod. Phys. 95, 011006 (2023).
@@
- [17] J. Ziegler, T. McJunkin, E. S. Joseph, S. S. Kalantre, B. Harpt, D. E. Savage, M. G. Lagally, M. A. Eriksson, J. M. Taylor, and J. P. Zwolak, Toward robust autotuning of noisy quantum dot devices, Phys. Rev. Appl. 17, 024069 (2022).
@@
- [18] J. P. Zwolak, T. McJunkin, S. S. Kalantre, J. Dodson, E. R. MacQuarrie, D. Savage, M. Lagally, S. Coppersmith, M. A. Eriksson, and J. M. Taylor, Autotuning of double-dot devices in situ with machine learning, Phys. Rev. Appl. 13, 034075 (2020).
@@
- [19] J. Ziegler, F. Luthi, M. Ramsey, F. Borjans, G. Zheng, and J. P. Zwolak, Tuning arrays with rays: Physics-informed tuning of quantum dot charge states, Phys. Rev. Appl. 20, 034067 (2023a).
@@
- [20] J. P. Zwolak, T. McJunkin, S. S. Kalantre, S. F. Neyens, E. R. MacQuarrie, M. A. Eriksson, and J. M. Taylor, Ray-based framework for state identification in quantum dot devices, PRX Quantum 2, 020335 (2021).
@@
- [21] F. Berritta, T. Rasmussen, J. A. Krzywda, J. van der Heijden, F. Fedele, S. Fallahi, G. C. Gardner, M. J. Manfra, E. van Nieuwenburg, J. Danon, A. Chatterjee, and F. Kuemmeth, Real-time two-axis control of a spin qubit, Nat. Commun. 15, 1676 (2024).
- [22] J. Hickie, B. van Straaten, F. Fedele, D. Jirovec, A. Ballabio, D. Chrastina, G. Isella, G. Katsaros, and N. Ares, Automated long-range compensation of an rf quantum dot sensor, Phys. Rev. Appl. 22, 064026 (2024).
@@
- [23] H. Moon, D. T. Lennon, J. Kirkpatrick, N. M. van Esbroeck, L. C. Camenzind, L. Yu, F. Vigneau, D. M. Zumbühl, G. A. D. Briggs, M. A. Osborne, D. Sejdinovic, E. A. Laird, and N. Ares, Machine learning enables completely automatic tuning of a quantum device faster than human experts, Nat. Commun. 11, 4161 (2020).
@@
- [24] G. J. Podd, S. J. Angus, D. A. Williams, and A. J. Ferguson, Charge sensing in intrinsic silicon quantum dot, Appl. Phys. Lett. 96, 082104 (2010).
@@
- [25] R. Durrer, B. Kratochwil, J. Koski, A. Landig, C. Reichl, W. Wegscheider, T. Ihn, and E. Greplova, Automated tuning of double quantum dots into specific charge states using neural networks, Phys. Rev. Appl. 13, 054019 (2020).
@@
- [26] J. Ziegler, F. Luthi, M. Ramsey, F. Borjans, G. Zheng, and J. P. Zwolak, Automated extraction of capacitive coupling for quantum dot systems, Phys. Rev. Appl. 19, 054077 (2023b).
@@
- [27] J. P. Zwolak, S. S. Kalantre, X. Wu, S. Ragole, and J. M. Taylor, QFlow lite dataset: A machine-learning approach to the charge states in quantum dot experiments, PLoS ONE 13, e0205844 (2018).
@@
- [28] F. Fedele, A. Chatterjee, S. Fallahi, G. C. Gardner, M. J. Manfra, and F. Kuemmeth, Simultaneous operations in a two-dimensional array of singlet-triplet qubits., PRX Quantum 2, 040306 (2021).
- [29] L. L. Sohn, L. P. Kouwenhoven, and G. Schön, Mesoscopic electron transport (Springer Netherlands, 1997).
@@
- [30] T. A. Baart, P. T. Eendebak, C. Reichl, W. Wegscheider, and L. M. K. Vandersypen, Computer-automated tuning of semiconductor double quantum dots into the single-electron regime, Appl. Phys. Lett. 108, 213104 (2016).
@@
- [31] We use a notation value(uncertainty) to express uncertainties, for example, . All uncertainties herein reflect the uncorrelated combination of single-standard deviation statistical and systematic uncertainties.